I. COMPUTATION

1. The momentum of a particle is 50 kg-m/s. If the velocity of the particle is 5 m/s, what is its mass? If the particle is located on a certain planet where the acceleration due to gravity is 5.87 m/s2, what is its weight?



1
Expert's answer
2022-06-06T14:58:15-0400

The mass of the particle

"m=\frac{p}{v}=\rm\frac{50\: kg\cdot m/s}{5\: m/s}=10\: kg"

The weight of the particle

"W=mg=\rm 10\: kg*5.87\: m/s^2=58.7\: N"